# Design of Pulse Shapes Based on Sampling with Gaussian Prefilter

### v3

@@ Abstract
Two new pulse shapes for communications are presented. The first pulse shape generates a set of pulses without intersymbol interference (ISI) or ISI-free for short. In the neighbourhood of the origin it is similar in shape to the classical cardinal sine function but is of exponential decay at infinity. This pulse shape is identical to the interpolating function of a recent sampling theorem with Gaussian prefilter. The second pulse shape is obtained from the first pulse shape by spectral factorization. Besides being also of exponential decay at infinity, it has a causal appearance since it is of superexponential decay for negative times. It is closely related to the orthonormal generating function considered earlier by Unser in the context of shift-invariant spaces. This pulse shape is not ISI-free but it generates a set of orthonormal pulses. The second pulse shape may also be used to define a receive matched filter so that at the filter output the ISI-free pulses of the first kind are recovered.
@@ I Introduction
The following notations and conventions are adopted: $L^{2}(\mathbb{R})$ is the space of square-integrable functions (or finite-energy signals) $f:\mathbb{R}\rightarrow\mathbb{C}\cup\{\infty\}$ with inner product $\langle f_{1},f_{2}\rangle=\int_{-\infty}^{\infty}f_{1}(x)\overline{f_{2}(x)}\,\mathrm{d}x.$ For the Fourier transform we use the definition $\hat{f}(\omega)=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}f(x)\,\mathrm{d}x$, where $x$ denotes time and $\omega$ angular frequency. $\ell^{2}(\mathbb{Z})$ is the space of square-summable, complex-valued sequences indexed by the integers. Finally, $\delta_{n}=0,\,n\in\mathbb{Z}\setminus\{0\}$, and $\delta_{0}=1$.
@@ II Sampling in Shift-Invariant Spaces and Localization Spaces Revisited
This theorem is due to Walter [4], who originally stated and proved it for orthonormal systems $\{\varphi(\cdot-n);n\in\mathbb{Z}\}$. The theorem was later generalized to Riesz bases by Unser [1], who also considered alternative generating functions for $V_{\lambda}(\varphi)$ like $\varphi_{\mathrm{int}}$ as above and $\varphi_{\mathrm{ortho}}$ (see below). In an attempt to retain the flavor of the original Whittaker–Kotelnikov–Shannon (WKS) sampling theorem [5], in [6] prior to sampling a prefilter
@@
where $\Lambda$ is as in (4). Note that because of the Riesz basis condition still imposed on $\varphi$ (see [6] for the full set of assumptions), which is equivalent to the existence of positive constants $A$ and $B$ (possibly depending on $\lambda$) so that
@@
(see [7] for a proof in the case of $\lambda=1$), the denominator in (7) never will vanish. We remark that in general $\Phi_{\mathrm{int}}\neq\boldsymbol{P}_{\varphi}\varphi_{\mathrm{int}}$. In the special case of a Gaussian prefilter function,
@@
where the parameter $\beta>0$ controls effective bandwidth, the following sampling theorem has been obtained [2], [6].
@@
For any $\lambda>0$ let the interpolating function $\Phi_{\mathrm{int}}$ be defined by (7). Then for any function $g\in\mathcal{P}_{\varphi}$ the function $\tilde{g}$ given by
@@
is again in $\mathcal{P}_{\varphi}$, it perfectly reconstructs $g$ at the sampling instants $x_{n}=n\lambda,\,n\in\mathbb{Z}$, and for all other $x\in\mathbb{R}$ the absolute approximation error $|g(x)-\tilde{g}(x)|$ becomes small as soon as $\lambda\leq 1/\beta$, then decaying superexponentially to zero as $\lambda\rightarrow 0$.
@@
$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\lambda}{\sqrt{2\pi}}\frac{\mathrm{e}^{-\frac{i\pi}{\tau}(\omega/\Lambda)^{2}}}{\sqrt{-i\tau}\vartheta_{3}(\omega/\Lambda,\tau)},$ (11)
@@
Note that in our context $\mathrm{i}\tau$ is always a negative real number and that $S_{0}(x)$ decays exponentially to zero as $x\rightarrow\pm\infty$.
@@ III Spectral Factorization of the Interpolating Function
The following identity of Euler holds true for any $q\in{\mathbb{C}},|q|<1,$ and $z\in\mathbb{C}$:
$1+\sum_{n=1}^{\infty}\frac{q^{\frac{1}{2}n(n-1)}}{(q;q)_{n}}z^{n}=\prod_{n=0}^{\infty}(1+zq^{n}).$ (16)
Jacobi’s triple product identity
@@
will also be needed. In our case always
@@
where the parameter $\tau$ is as in (13). The special $q$-Pochhammer symbols
@@
is the Gaussian function (9) normalized to unit energy, i.e. $\int_{\mathbb{R}}|\phi(x)|^{2}\,\mathrm{d}x=1$, it holds that
@@
Case $n=-1,-2,\ldots:$ After substitution in (22) of $x$ by $1-x$ we obtain
@@
where integration in the contour integral is performed counterclockwise around the unit circle. Since the integrand function is analytic within a neighbourhood of the closed unit disc, we obtain by means of Cauchy’s integral theorem that $a_{n}=0,\,n=-1,-2,\ldots$
@@
It can be shown that $V_{\lambda}(\varphi_{\mathrm{ortho}})=V_{\lambda}(\varphi)$, where $V_{\lambda}(\varphi)$ is again the shift-invariant space (2) with Gaussian generating function $\varphi$ as given in (9). In Fig. 2, the function $\varphi_{\mathrm{ortho}}$ together with some of its translates is depicted for bandwidth parameter $\beta=100$ and $\lambda=3/\beta$.
@@ IV-A ISI-Free Pulses
(the left-hand equation being Poisson’s summation formula; e.g., [1]), which is equivalent to $\Phi_{\mathrm{int}}(n\lambda)=\delta_{n},\,n\in\mathbb{Z}$. Therefore, the set of shifted pulses $\{\Phi_{\mathrm{int}}(x-n\lambda);n\in\mathbb{Z}\}$ is ISI-free at points in time $x_{n}=n\lambda$. The right-hand equation in (23) is, of course, the well-known Nyquist criterion for ISI-free pulses [10]. Concerning the use of the also ISI-free pulses generated by (15) (for arbitrary $\beta>0,\,\lambda>0$) refer to [11].
@@ IV-B1 Orthonormal Pulses
It is well-known [1] that for an arbitrary function $\varphi\in L^{2}(\mathbb{R})$ the system of functions $\{\varphi(x-n\lambda);n\in\mathbb{Z}\}$ forms an orthonomal system in $L^{2}(\mathbb{R})$ if and only if in (8) $A=B=1$ may be chosen. For the function $\varphi_{\mathrm{ortho}}$ of Theorem 3 this is true due to Eq. (21) and the right-hand equation in (23). Therefore, $\varphi_{\mathrm{ortho}}$ may be used as a pulse shape to generate a set $\{\varphi_{\mathrm{ortho}}(x-n\lambda);\,n\in\mathbb{Z}\}$ of orthonormal pulses.
@@ IV-B2 Matched Filter
At the receiver, the filter $\boldsymbol{P}_{\varphi_{\mathrm{ortho}}}$ obtained by replacing the prefilter function $\varphi$ in (5) with $\varphi_{\mathrm{ortho}}$ forms a matched filter allowing optimal detection of the pulses $\varphi_{\mathrm{ortho}}(x-n\lambda)$ at points in time $x_{n}=n\lambda,\;n\in\mathbb{Z},$ in the presence of noise. Thanks to orthogonality, overlap of consecutive pulses is irrelevant. Moreover, since [cf. (6) and Eq. (21)]
@@
In the light of the last application, the two proposed pulse shapes $\Phi_{\mathrm{int}}$ and $\varphi_{\mathrm{ortho}}$ are seen to correspond to the classical pulse shapes produced by a raised-cosine filter or a root raised-cosine filter respectively [10]. Recall that the proposed pulse shapes decay (super-)exponentially as $x\rightarrow\pm\infty$ whereas the classical pulse shapes merely decay as $x^{-3}$.
@@ acknowledgement
The author wishes to thank Ayush Bhandari for interesting discussions and helpful suggestions concerning some topics of this paper.

### v4

@@ Abstract
Two new pulse shapes for communications are presented. The first pulse shape generates a set of pulses without intersymbol interference (ISI) or ISI-free for short. In the neighborhood of the origin it is similar in shape to the classical cardinal sine function but is of exponential decay at infinity. This pulse shape is identical to the interpolating function of a generalized sampling theorem with Gaussian prefilter. The second pulse shape is obtained from the first pulse shape by spectral factorization. Besides being also of exponential decay at infinity, it has a causal appearance since it is of superexponential decay for negative times. It is closely related to the orthonormal generating function considered earlier by Unser in the context of shift-invariant spaces. This pulse shape is not ISI-free but it generates a set of orthonormal pulses. The second pulse shape may also be used to define a receive matched filter so that at the filter output the ISI-free pulses of the first kind are recovered.
@@ I Introduction
The following notations and conventions are adopted: $L^{2}(\mathbb{R})$ is the space of square-integrable functions (or finite-energy signals) $f:\mathbb{R}\rightarrow\mathbb{C}\cup\{\infty\}$ with inner product $\langle f_{1},f_{2}\rangle=\int_{-\infty}^{\infty}f_{1}(x)\overline{f_{2}(x)}\,\mathrm{d}x$ and norm $\|f\|=\langle f,f\rangle^{1/2}$. For the Fourier transform we use the definition $\hat{f}(\omega)=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}f(x)\,\mathrm{d}x$, where $x$ denotes time and $\omega$ angular frequency. $\ell^{2}(\mathbb{Z})$ is the space of square-summable, complex-valued sequences indexed by the integers. Finally, $\delta_{n}=0,\,n\in\mathbb{Z}\setminus\{0\}$, and $\delta_{0}=1$.
@@ II Sampling in Shift-Invariant Spaces and Localization Spaces Revisited
This theorem is due to Walter [4], who originally proved it for orthonormal bases $\{\varphi(\cdot-n);n\in\mathbb{Z}\}$. The theorem was later extended to Riesz bases by Unser [1], who also considered alternative generating functions for $V_{\lambda}(\varphi)$ like $\varphi_{\mathrm{int}}$ as above and $\varphi_{\mathrm{ortho}}$ (see below). In an attempt to retain the flavor of the original Whittaker–Kotelnikov–Shannon (WKS) sampling theorem [5], in [6] prior to sampling a prefilter
@@
where $\Lambda$ is as in (4). Note that because of the Riesz basis condition still imposed on $\varphi$ (see [6] for the full set of assumptions), which is equivalent to the existence of positive constants $A$ and $B$ (possibly depending on $\lambda$) so that [7]
@@
the denominator in (7) never will vanish. We remark that in general $\Phi_{\mathrm{int}}\neq\boldsymbol{P}_{\varphi}\varphi_{\mathrm{int}}$. In the special case of a Gaussian prefilter function,
@@
where the parameter $\beta>0$ controls effective bandwidth, the following generalized sampling theorem has been obtained [2], [6].
@@
For any $\lambda>0$ let the interpolating function $\Phi_{\mathrm{int}}$ be defined by (7). Then for any function $g\in\mathcal{P}_{\varphi}$, where $g=\boldsymbol{P}_{\varphi}f,\,f\in L^{2}(\mathbb{R})$, the function $\tilde{g}$ given by
@@
is again in $\mathcal{P}_{\varphi}$, it perfectly reconstructs $g$ at the sampling instants $x_{n}=n\lambda,\,n\in\mathbb{Z}$, and for all other $x\in\mathbb{R}$ the squared relative error $\left(|g(x)-\tilde{g}(x)|/\|f\|\right)^{2}$ becomes small as soon as $\lambda\leq 1/\beta$, then decaying superexponentially to zero as $\lambda\rightarrow 0$.
@@
$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\lambda}{\sqrt{2\pi}}\frac{\mathrm{e}^{-\frac{\mathrm{i}\pi}{\tau}(\omega/\Lambda)^{2}}}{\sqrt{-\mathrm{i}\tau}\vartheta_{3}(\omega/\Lambda,\tau)},$ (11)
@@
Note that in our context $\mathrm{i}\tau$ is always a negative real number and that $\Phi_{\mathrm{int}}(x)$ decays exponentially to zero as $x\rightarrow\pm\infty$.
@@ III Spectral Factorization of the Interpolating Function
The following identities of Euler hold true for $q\in\mathbb{C},|q|<1$:
$1+\sum_{n=1}^{\infty}\frac{q^{\frac{1}{2}n(n-1)}}{(q;q)_{n}}z^{n}=\prod_{n=0}^{\infty}(1+zq^{n}),\,z\in\mathbb{C},$ (16)
$1+\sum_{n=1}^{\infty}\frac{z^{n}}{(q;q)_{n}}=\prod_{n=0}^{\infty}(1-zq^{n})^{-1},\,z\in\mathbb{C},|z|<1.$ (17)
Only Eq. (16) will be needed in the proof of the next theorem, where Jacobi’s triple product identity
@@
will play a central role. In our case always
@@
where parameter $\tau$ is as in (13). The special $q$-Pochhammer symbols
@@
is the Gaussian function $\varphi$ in (9) normalized to unit energy, i.e., $\int_{\mathbb{R}}|\phi(x)|^{2}\,\mathrm{d}x=1$, it holds that
@@
Case $n=-1,-2,\ldots:$ After substitution in (23) of $x$ by 1x1-x we obtain
@@
where integration in the contour integral is performed counterclockwise around the unit circle. Since the integrand function is analytic within a neighborhood of the closed unit disc, we obtain by means of Cauchy’s integral theorem that $a_{n}=0,\,n=-1,-2,\ldots$
@@
In Fig. 2, the function $\varphi_{\mathrm{ortho}}$ along with some of its translates is depicted for bandwidth parameter $\beta=100$ and $\lambda=3/\beta$.
Proposition 1
The function $\phi$ in (21) has for $x\in\mathbb{R}$ the representation
$\phi(x)=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{{(q^{2};q^{2})_{n}}}\,\varphi_{\mathrm{ortho}}(x-n\lambda),$ (24)
where $q$ is as in Theorem 3
Proof:
The function $\varphi_{\mathrm{ortho}}$ as given in Eq. (20) has the Fourier transform
$\hat{\varphi}_{\mathrm{ortho}}(\omega)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}\varphi_{\mathrm{ortho}}(x)\,\mathrm{d}x$
$=\frac{1}{Q_{0}^{1/2}}\sum_{n=0}^{\infty}\frac{(-q)^{n}}{(q^{2};q^{2})_{n}}\,\mathrm{e}^{-\mathrm{i}n\lambda\omega}\hat{\phi}(\omega)$
$=\frac{1}{Q_{0}^{1/2}}\sum_{n=0}^{\infty}\frac{(-q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}}{(q^{2};q^{2})_{n}}\,\hat{\phi}(\omega).$
Because of Eq. (17) it holds that
$\sum_{n=0}^{\infty}\frac{(-q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}}{(q^{2};q^{2})_{n}}=\prod_{n=0}^{\infty}(1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}\cdot q^{2n})^{-1}.$
We thus obtain
$\hat{\varphi}_{\mathrm{ortho}}=\frac{1}{Q_{0}^{1/2}}\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]^{-1}\hat{\phi}(\omega),$
which is equivalent to
$\hat{\phi}(\omega)=Q_{0}^{1/2}\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]\,\hat{\varphi}_{\mathrm{ortho}}(\omega).$ (25)
Because of Eq. (16) it furthermore holds that
$\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]$ $=\sum_{n=0}^{\infty}\frac{q^{n(n-1)}}{(q^{2};q^{2})_{n}}(q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}$
$=\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{(q^{2};q^{2})_{n}}\,\mathrm{e}^{-\mathrm{i}n\lambda\omega}.$
Inserting the last expression into Eq. (25) and inverting the Fourier transform yields
$\phi(x)$ $=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{(q^{2};q^{2})_{n}}\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}\omega(x-n\lambda)}\hat{\varphi}_{\mathrm{ortho}}(\omega)\,\mathrm{d}\omega$
$=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{{(q^{2};q^{2})_{n}}}\,\varphi_{\mathrm{ortho}}(x-n\lambda),$
which concludes the proof of Proposition 1. ∎
Corollary 1
Let the Gaussian generator $\varphi$ be as given in (9) and let the function $\varphi_{\mathrm{ortho}}$ be as given in Eq. (20). Then for any $\lambda>0$ it holds that
$V_{\lambda}(\varphi_{\mathrm{ortho}})=V_{\lambda}(\varphi).$ (26)
Proof:
For any generating function $\varphi\in L^{2}(\mathbb{R})$ and any $\lambda>0$, the shift-invariant space $V_{\lambda}(\varphi)$ as given in (2) may also be defined as the closed linear span of the system of functions $\{\varphi(\cdot-n\lambda);\,n\in\mathbb{Z}\}$ in $L^{2}(\mathbb{R})$; in particular, $V_{\lambda}(\varphi)$ is a closed linear subspace of $L^{2}(\mathbb{R})$. In the case of the above Gaussian generator $\varphi$, due to Theorem 3, $\varphi_{\mathrm{ortho}}\in V_{\lambda}(\phi)=V_{\lambda}(\varphi)$ so that
$V_{\lambda}(\varphi_{\mathrm{ortho}})\subseteq V_{\lambda}(\varphi).$ (27)
On the other hand, by reason of Proposition 1, $\varphi\propto\phi\in V_{\lambda}(\varphi_{\mathrm{ortho}})$ so that
$V_{\lambda}(\varphi)\subseteq V_{\lambda}(\varphi_{\mathrm{ortho}}).$ (28)
The set inclusion (27) in combination with the set inclusion (28) results in the set equality (26), which concludes the proof of Corollary 1. ∎
@@ IV-A ISI-Free Pulses
(the first equation being Poisson’s summation formula; e.g., [1]), which is equivalent to $\Phi_{\mathrm{int}}(n\lambda)=\delta_{n},\,n\in\mathbb{Z}$. Therefore, the set of shifted pulses $\{\Phi_{\mathrm{int}}(x-n\lambda);n\in\mathbb{Z}\}$ is ISI-free at points in time $x_{n}=n\lambda$. The second equation in (29) is, of course, an instance of the well-known Nyquist criterion for ISI-free pulses [10]. Concerning the use of the also ISI-free pulses generated by (15) (for arbitrary positive parameters $\beta$ and $\lambda$) refer to [11].
@@ IV-B1 Orthonormal Pulses
For any function $\varphi\in L^{2}(\mathbb{R})$ the system of functions $\{\varphi(x-n\lambda);n\in\mathbb{Z}\}$ forms an orthonormal system in $L^{2}(\mathbb{R})$ if and only if in (8) $A=B=1$ may be chosen [1]. For the function $\varphi_{\mathrm{ortho}}$ of Theorem 3 this is true because of Eq. (22) and the second equation in (29). Therefore, $\varphi_{\mathrm{ortho}}$ may be used as a pulse shape to generate a set $\{\varphi_{\mathrm{ortho}}(x-n\lambda);\,n\in\mathbb{Z}\}$ of orthonormal pulses.
@@ IV-B2 Matched Filter
At the receiver, the filter $\boldsymbol{P}_{\varphi_{\mathrm{ortho}}}$ obtained by replacing the prefilter function $\varphi$ in (5) with $\varphi_{\mathrm{ortho}}$ forms a matched filter allowing optimal detection of the pulses $\varphi_{\mathrm{ortho}}(x-n\lambda)$ at points in time $x_{n}=n\lambda,\;n\in\mathbb{Z},$ in the presence of noise. Due to orthogonality, overlap of adjacent pulses is of no relevance. Moreover, since it holds that [cf. (6) and Eq. (22)]
@@
In the light of the last application, the two proposed pulse shapes $\Phi_{\mathrm{int}}$ and $\varphi_{\mathrm{ortho}}$ are seen to correspond to the classical pulse shapes produced by a raised-cosine filter or a root raised-cosine filter respectively [10]. Recall that the proposed pulse shapes decay (super-)exponentially to zero as $x\rightarrow\pm\infty$ whereas the classical pulse shapes merely decay order of $x^{-3}$.
@@ acknowledgement
The author wishes to thank Ayush Bhandari for interesting discussions and helpful suggestions concerning an earlier draft of this paper.